# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v1

@@ 6.2 Scenery Tag Test
Table 1: Metric results on Scenery Tag Test.
As shown in Table 1, applying TIPO significantly improves FDD performance and AI Corrupt score. The improvement in FDD score suggests that TIPO-generated prompts lead to images that more closely align with the original dataset distribution. Furthermore, the enhanced AI Corrupt score indicates that TIPO aids in generating more coherent images, potentially reducing corruption or distortion in the output.
The results from the direct LLM approach (GPT4o-mini) exemplify a key concern in prompt engineering: while achieving high aesthetic scores, it performs poorly in terms of FDD. This discrepancy underscores the potential pitfall of using LLMs for prompt generation without considering dataset alignment. Such models may produce well-constructed prompts that yield visually appealing images but fail to capture the nuanced distribution of the text-to-image dataset. In contrast, TIPO demonstrates a balance between aesthetic quality and dataset fidelity, offering a more robust solution for prompt optimization in text-to-image generation tasks.
6.3 Short/Truncated Long Test
This test evaluates TIPO’s performance on enhancing natural language prompts which are already contains some useful info for model to generate reasonable images. We randomly selected 10k entries each from [9] and [25]. For each entry, we extracted the short caption and the truncated long caption (limited to a maximum of two sentences). These captions serve as user input prompts for our test. Using these prompts, we generated one image per prompt with the SDXL-1.0-base [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 model [37].
@@ 6.3 Short/Truncated Long Test
Table 2: Metric result in Short/Truncated-long test of prompt generation
Table 2 presents the results of our Short/Truncated-long test for prompt generation. Unlike the scenery tag test (Section 6.2), we observe that the original prompts in this test already produce decent images. Interestingly, all four methods result in worse FDD scores compared to the original prompts, suggesting that the prompt generators introduced new elements not present in the selected 10,000 entries. Given the size of the training set, this outcome is not entirely unexpected.
Examining the Aesthetic and AI Corrupt Scores, we notice that the Direct LLM approach achieves the highest aesthetic score, mirroring the results from the scenery tag test. However, its AI Corrupt Score and FDD value indicate potential issues with OOD inputs, which may occasionally hinder the model’s ability to generate coherent images from its prompts.
TIPO, on the other hand, demonstrates a more balanced performance. It achieves the best AI Corrupt Score, the second-best Aesthetic Score, and a reasonable FDD value. This balanced profile suggests that TIPO is the most suitable choice for real-world applications of modern text-to-image models, where a combination of aesthetic appeal, image coherence, and distribution alignment is crucial.
@@ 7 Conclusion
In this paper, we introduced TIPO (Text to Image with text Presampling for Prompt Optimization), a novel framework designed to enhance the quality of Text-to-Image (T2I) models through automatic prompt engineering. By leveraging Large Language Models (LLMs) to extend and refine user-provided prompts, TIPO effectively bridges the gap between simple user inputs and the detailed, content-rich prompts required for high-quality image generation.
Our comprehensive evaluation demonstrates TIPO’s effectiveness across multiple dimensions of image quality and fidelity. The framework consistently produced prompts that led to images closely aligned with original dataset distributions, while simultaneously improving aesthetic appeal and reducing the likelihood of corrupted or nonsensical outputs. This balanced performance underscores TIPO’s robustness and suitability for real-world applications of T2I models, where a combination of distribution alignment, visual appeal, and image coherence is crucial.
Notably, our experiments highlight the critical role of prompt engineering in advancing T2I technology. The improvements achieved through TIPO’s prompt modifications often rivaled or surpassed performance differences between various T2I model architectures. This finding emphasizes that effective prompt engineering can significantly enhance model performance without necessitating architectural changes, potentially offering a more efficient path to improving T2I systems.
The success of TIPO opens several promising avenues for future research:
  1. 1.
    Broader Application: Exploring TIPO’s applicability to other generative tasks, such as text generation or audio synthesis, could further demonstrate the versatility and impact of automatic prompt engineering.
  2. 2.
    Interactive Systems: Incorporating TIPO into interactive applications where users can iteratively refine prompts in real-time may enhance user experience and enable more precise control over generated content.
  3. 3.
    Advanced Alignment Techniques: Investigating more sophisticated methods for aligning user inputs with dataset distributions could further enhance the performance and reliability of generative models.
  4. 4.
    Cross-Modal Applications: Adapting TIPO’s principles to facilitate cross-modal generation tasks, such as text-to-video or image-to-text, could expand its utility in multimodal AI systems.
In conclusion, TIPO represents a significant advancement in the field of text-to-image generation, offering a versatile and efficient solution for prompt optimization. By automating the refinement and extension of user prompts, TIPO not only improves the quality and specificity of generated images but also highlights the importance of aligning user inputs with underlying data distributions. As generative models continue to evolve, frameworks like TIPO will be instrumental in unlocking their full potential, ensuring they meet diverse user needs with precision and creativity while pushing the boundaries of what’s possible in AI-generated content.
@@ A.3 Coyo HD 11M
The Coyo HD 11M dataset [9] consists of 11.4 million high-density, high-definition images, with 22.8 million synthetic captions generated from the Coyo-700M dataset. The ”HD” refers to both the high resolution of images (filtered to have at least 512 pixels on the shortest edge) and the high concept density, addressing the issues found in previous alt-text image pair datasets, which often included low-quality or low-concept images. Synthetic captions for this dataset were produced using the LLaVA-Next-8B model [30] built upon LLaMA 3 [3] with post-processing for conciseness and clarity. This dataset provides high-quality visual and textual pairs, ideal for training advanced vision-language models.
@@
Appendix B Implementation Detail
In this appendix, we provide all the missing details about our dataset construction process that are not mentioned in Section 4.4 and 5.2.
B.1 Training Settings
TIPO-200M stage1 TIPO-200M stage2 TIPO-500M
@@ B.1 Training Settings
Type Pretrain Finetune Pretrain
Vocab Size 32013 - 32013
Hidden Dim 768 - 1280
Attention Heads 12 - 20
MLP Dim 2304 - 3840
Hidden Layers 20 - 20
Model Parameters 203M - 508M
Max Learning Rate 2e-4 5e-5 2e-4
Optimizer AdamW AdamW AdamW
betas 0.9, 0.98 0.9, 0.98 0.9, 0.98
weight decay 0.01 0.01 0.01
Dataset GBC, Dan Coyo, GBC, Dan Coyo, GBC, Dan
max context length 512 1024 1024
global batch size 2048 2048 3584
Token Seen 22.625B 18.339B 31.274B
Hardware 4 x RTX3090 4 x RTX3090 8 x H100
Training Time (wall) 150 hour 270 hour 100 hour
Table 3: Training settings and details for TIPO-200M and TIPO-500M.
B.2 Inference Pipeline
@@
Figure 5: The inference workflow in TIPO, demonstrating how different task combinations are used to refine or expand tag-based and natural language prompts.
@@ Appendix C Evaluation Statistics
C.1 Scenery Tag Test
@@
Figure 6: The box plot for scenery tag test.
The box plot shown in Figure 6 presents the aesthetic scores and AI corruption scores for results from the scenery tag test described in Section 6.2. The analysis clearly demonstrates that TIPO significantly outperforms all other methods, with a substantial margin of improvement.
C.2 Short/Truncated Long Test
Figures 7 and 8 display the box plots of aesthetic scores and AI corruption scores obtained from the short/truncated long test detailed in Section E.2. While the box plots reveal that the differences in performance between various methods are subtle, the AI corruption scores provide valuable insights. Specifically, these scores indicate that implementations supported by TIPO produce more stable output images compared to other methods.
(a) The box plot for the Aesthetic Score result of short caption input.
(b) The box plot for the AI Corrupt score result of short caption input.
Figure 7: The box plot for short caption input.
(a) The box plot for the Aesthetic Score result of truncated long caption input.
(b) The box plot for the AI Corrupt score result of truncated long caption input.
Figure 8: The box plot for truncated long caption input.
@@ Appendix E Image Examples
E.1 Scenery Tag Test
(a) Scenery Tag Only
(b) TIPO + Scenery Tag
@@ E.1 Scenery Tag Test
E.2 Short/Truncated Long Test
@@
(c) Truncated Long Caption
(d) TIPO-Extended Caption

### v2

@@ 5.2 In-domain test regard to scenery tag
Table 1: Metric results on in-domain test. Bold values indicate the best performance, and underlined values indicate the second-best. Original refers to images generated using the original text captions; LLM uses GPT-4o-mini [2]; the remaining columns represent results from Prompt DB [25], Promptist [26], and TIPO, respectively.
As shown in Table 1, TIPO-generated prompts improve FDD scores by aligning images more closely with the original dataset distribution. Enhanced AI Corrupt scores show that TIPO reduces image corruption, creating more coherent outputs by replacing vague descriptions (e.g., detailed background) with specific elements (e.g., garden filled with forget-me-nots and blue butterflies). The direct LLM approach highlights a key issue in prompt engineering: it achieves high aesthetic scores but compromises FDD, reducing the diversity of images.
We provide more comparative visual result in Figure 5. The results reveal significant biases in Prompt DB  [25] and RL-based [26] approaches. PromptDB and Direct-LLM method exhibits notable bias, by generating mountain/lake scene majorly. Meanwhile, Promptist produces diverse near-future fantasy sceneries, however often produces corrupted images due to unreasonable vocabulary usage (4K, masterpiece, instagram), supposedly introduced by user prompts.
By contrast, TIPO consistently reproduces a balanced output distribution, incorporating both natural and architectural elements, effectively mitigating the biases observed in other methods.
(a) Original prompt best aesthetic
(b) GPT-4o-mini best aesthetic
(c) TIPO best aesthetic
Figure 7: Top-10% aesthetic score images generated by the SD-3.5-Large model [19]. Rows show samples from Original, Direct LLM-Modified prompts, and TIPO-extended prompts. Direct LLM images focus primarily on natural scenery, while generated image from TIPO-extended prompt display a wider variety, including urban scenes and human figures.
5.3 In-domain prompt generation test
We aim to test each models’ performance on regard to reconstructing / upsampling the given prompt, in given dataset distribution. By comparing original ground truth and generated outputs, we can directly compare FDD, aesthetic score, AI corrupt scores.
We randomly selected 10k entries each from [12] and [29]. For each entry, we extracted the short caption and the truncated long caption (limited to a maximum of two sentences). These captions serve as user input prompts $p_{s}$ for in-domain prompt generation test. Using these prompts, we generated single image per prompt with the SDXL-1.0-base [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 model [48].
@@ 5.3 In-domain prompt generation test
Table 2: Performance comparison in the in-domain prompt generation test. Bold values indicate the best performance, and underlined values indicate the second-best. Direct LLM achieves the highest Aesthetic Score but struggles with OOD issues. TIPO provides a balanced outcome, excelling in AI Corrupt Score and maintaining strong alignment, making it ideal for practical applications.
Table 2 presents the results of our Short/Truncated caption test for prompt generation. Compared to in-domain test, Section 5.2, which might be considered slightly out of domain by SDXL model, we observed that prompts covered by this test produces meaningful, coherent images with sufficient detail level. FDD scores were degraded compared to original-ground truth prompts, not truncated long caption-which implies prompt extension functions resulted lack of diversity, or unmatched components in generation.
We observe that Direct LLM approach achieves the highest aesthetic scores by wide vocabulary with user preference matching, however results low FDD, AI corruption score, which implies unmatched vocabularies are being used in-prompt.
Original LLM Prompt DB Promptist TIPO(Ours)
Top 10% $\uparrow$ 10.98 7.84 12.17 10.12 17.33
Worst 10% $\uparrow$ 12.10 9.76 15.23 14.14 13.92
Table 3: Vendi Scores [23] for the top and worst 10% aesthetic images across prompt generation methods. Higher value indicate higher diversity of samples. TIPO achieves high score in the top 10% (17.33), showing greater visual diversity as well as worst 10%. Result suggests TIPO’s balanced and robust variety in generated images. Meanwhile, LLM has the lowest scores in both categories, indicating limited diversity and repetition.
5.4 Out of focus evaluation.
In this section, we analyze TIPO’s behavior in out-of-focus model, SD-3.5-Large [^4] https://huggingface.co/stabilityai/stable-diffusion-3.5-large built upon MMDiT architecture [21]. We compose original tag, natural language prompt created by GPT-4o-mini, tag+natural language extended prompt by TIPO (ours), then feed to SD-3.5-Large inference pipeline.
Analysis of Vendi Scores [23] (Table 3) reveals contrasting patterns: Direct LLM-generated images show limited diversity despite high aesthetic scores, while TIPO maintains conceptual diversity across all quality levels. However, TIPO’s training on diverse datasets [45, 65, 12, 29] introduces domain-specific elements that may not align with SD3.5L’s training distribution. This mismatch demonstrates our core finding: prompt generation models perform optimally when trained on domain-matched prompt corpora.
The balanced profile of TIPO suggests that TIPO can work as suitable solution for prompt upsampling task, which can accurately project trained prompt with few initial word input.
@@ 6 Conclusion
In this paper, we introduced TIPO (Text to Image with text Presampling for Prompt Optimization), a novel framework designed to enhance the quality of Text-to-Image (T2I) models through automatic prompt engineering. TIPO represents a significant advancement in the field of T2I generation, offering a versatile and efficient solution for prompt optimization. By automating the refinement and extension of user prompts. We also highlights the importance of aligning user inputs with underlying data distributions through the comparisons between different methods for prompt engineering. As generative models continue to evolve, frameworks like TIPO will be instrumental in unlocking their full potential, ensuring they meet diverse user needs with precision and creativity while pushing the boundaries of what’s possible in generative models.
@@ A.3 Coyo HD 11M
The Coyo HD 11M dataset [12] consists of 11.4 million high-density, high-definition images, with 22.8 million synthetic captions generated from the Coyo-700M dataset. The “HD” refers to both the high resolution of images (filtered to have at least 512 pixels on the shortest edge) and the high concept density, addressing the issues found in previous alt-text image pair datasets, which often included low-quality or low-concept images. Synthetic captions for this dataset were produced using the LLaVA-Next-8B model [37] built upon LLaMA 3 [3] with post-processing for conciseness and clarity. This dataset provides high-quality visual and textual pairs, ideal for training advanced vision-language models.
@@
A.4 Stable Diffusion XL
The Stable Diffusion XL [48] improves upon earlier versions [54] with a significantly larger UNet backbone and dual text encoders, CLIP ViT-L [49] and OpenCLIP ViT-bigG [30], to enhance text conditioning. It supports resolutions up to 1024×1024 and is designed for diverse image generation. Although explicit dataset details are not provided, figures in the paper confirm that SDXL accepts both natural language (NL) prompts and tags as input.
A.5 Stable Diffusion 3.5
The Stable Diffusion 3.5 (SD-3.5) incorporates a cutting-edge architecture, MMDiT [21], along with the Rectified Flow [38, 4, 35] formulation for improved text-to-image generation. The model utilizes triple text encoders—CLIP/ViT-L, OpenCLIP/ViT-G, and T5-XXL [50]—for robust text conditioning. SD-3.5 supports resolutions up to 1024×1024 and employs a 50/50 mix of original and CogVLM-generated captions for training. While explicit dataset details are not provided, figures in the paper confirm its capability to process both natural language prompts and tags.
In this study, we use the SD-3.5-Large [^8] https://huggingface.co/stabilityai/stable-diffusion-3.5-large (8B parameters) with FP8 inference on RTX 3090 or RTX 4090 GPUs.
Appendix B Implementation Details
In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline and the model’s property which are not mentioned in Section 4.3 and 4.4.
B.1 Training Data Construction
This section details our methodology for constructing and preprocessing training data to ensure robust model performance across various input scenarios.
Length Control
To systematically control output prompt length, we implement a structured length categorization system using special length tags. These tags enforce specific constraints on both tag counts and natural language sentence lengths. For instance, the <long> tag specifies that the corresponding prompt must contain between 36 and 52 tags (inclusive), accompanied by 4 to 8 sentences of natural language description. We define four distinct length categories, each with strict bounds for both tag count and sentence length.
Type Very Short Short Long Very Long
Tags (count) 18 36 48 72
NL (sentences) 2 4 8 18
Table 4: Maximum length specifications for each category and caption type. For each category, the actual count/length must not exceed these values.
Random Augmentation
To enhance input diversity and better simulate real-world usage patterns, we implement several data augmentation strategies:
  • Metadata Tags: For tags representing image metadata (e.g., artist, character, aspect ratio), we employ two randomization techniques:
    • Random removal of metadata tags
    • Random repositioning of metadata tags to the end of the prompt, after all content-related descriptions
    This approach encourages the model to handle varying metadata positions and availability, while maintaining the ability to infer metadata relationships from the content descriptions.
  • Content Tags: For tags describing image content (e.g., objects, actions, attributes), we implement:
    • Random shuffling of tag order within the content section
    • Length-based truncation to meet target length constraints while preserving key content information
  • Natural Language: For natural language descriptions exceeding length limitations, we employ selective sentence removal, targeting middle sentences to preserve both context-setting opening sentences and concluding details. This maintains coherent narrative flow while meeting target length requirements.
These augmentation strategies create a more diverse training dataset that better reflects real-world prompt variations, improving the model’s robustness and adaptability to different input styles and formats.
B.2 Training Settings and Model Configurations
TIPO-100M TIPO-200M stage1 TIPO-200M stage2 TIPO-500M
@@ B.2 Training Settings and Model Configurations
Type Pretrain Pretrain Finetune Pretrain
Vocab Size 32013
Hidden Dim 640 768 - 1280
Attention Heads 10 12 - 20
MLP Dim 2240 2304 - 3840
Hidden Layers 10 20 - 20
Model Parameters 100M 203M - 508M
Max Learning Rate 5e-4 2e-4 5e-5 2e-4
Optimizer AdamW
LR scheduler Cosine Annealing LR
betas 0.9, 0.98
weight decay 0.01
Dataset Coyo, GBC, Dan GBC, Dan Coyo, GBC, Dan Coyo, GBC, Dan
Epoch 1 5 3 5
max context length 512 512 1024 1024
global batch size 1024 2048 2048 3584
Token Seen 6.0240B 22.625B 18.339B 31.274B
Hardware 4 x RTX3090 4 x RTX3090 4 x RTX3090 8 x H100
Training Time (wall) 22.5 hour 150 hour 270 hour 100 hour
Table 5: Training settings and details for TIPO-100M, 200M, and 500M. The datasets include CoyoHD11M (Coyo), GBC10M (GBC), and Danbooru2023 (Dan). Stage 2 additionally incorporates Pixtral [1] to generate NL captions from Danboodu2023 dataset.
Tokenizer and Task Tokens
TIPO employs a vocabulary derived from LLaMA2 [62] consisting of 32,000 tokens, with additional tokens (13 tokens) specifically designated for task and length control or placeholders. This extended vocabulary includes task identifiers and length modifiers to ensure flexibility across different prompt types:
  • Placeholder Token (1 token):
        <|empty|>
  • Task Tokens (8 tokens):
        <|gen_meta|>, <|tag_to_long|>, <|short_to_tag|>,
        <|long_to_tag|>, <|short_to_long|>, <|short_to_tag_to_long|>,
        <|short_to_long_to_tag|>, <|tag_to_short_to_long|>
  • Length Tokens (4 tokens):
        <|very_short|>, <|short|>, <|long|>, <|very_long|>
Optimizer and Learning Schedule
Training is performed using the AdamW optimizer [39], with a cosine annealing learning rate scheduler [40]. The optimizer parameters include $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01. Maximum learning rates are adjusted per model size, as outlined in Table 5.
Training Configurations
TIPO models are trained in multiple stages. Table 5 summarizes the configurations for pretraining and fine-tuning TIPO-100M, TIPO-200M, and TIPO-500M. Both pretraining and finetuning was conducted on datasets like Danbooru2023 [45], GBC10M [29], and CoyoHD11M [12].
Augmented Task Representation
Each dataset entry undergoes random task assignment and splitting to simulate a wide range of input-output mappings, effectively increasing the dataset size. For example, a single entry may contribute to tasks like short_to_tag or tag_to_long, with length modifiers dynamically controlling the output verbosity. This approach ensures the model can handle diverse tasks while maintaining robust generalization.
Hardware and Time Requirements
Training was conducted on NVIDIA RTX3090 GPUs for smaller models and H100 GPUs for TIPO-500M. Total wall-clock training times ranged from 22.5 hours for TIPO-100M to 270 hours for fine-tuning TIPO-200M.
Token Seen and Effective Training
Non-padding tokens are used to measure the effective token count during training, ensuring efficiency given the short and variable data lengths. Table 5 details the total tokens seen per model and training stage, illustrating the comprehensive exposure to diverse data entries.
[…]
Figure 8: TIPO inference workflow, with solid arrows denoting the primary generation steps and dashed arrows indicating alternative generation paths within the same cycle. ¡TOKEN¿— represents special tokens, with all tokens detailed in Section B.2.
B.3 Inference Pipeline
@@
B.4 Model Speed Comparison
We conducted comprehensive speed tests of our 100M, 200M and 500M parameter models using the inference pipeline described in Section 5.2. Each prompt requires two sequential generation steps. Our primary metric is average tokens generated per second, which reflects real-world task performance rather than theoretical maximum throughput.
The evaluation was performed using llama.cpp [24], an efficient C++ implementation that provides optimized support for various hardware accelerators including CUDA, HIP, and Apple Metal.
TIPO-100M TIPO-200M TIPO-500M
Hardware Platform tok/sec gen time tok/sec gen time tok/sec gen time
M1 Max (32 GPU cores) 339.4 0.66 190.0 1.23 119.4 2.02
RTX 3090 558.5 0.42 341.4 0.69 289.8 0.81
RTX 4090 742.9 0.29 454.5 0.51 359.7 0.63
Table 6: Model performance comparison across different hardware platforms. Tokens per second (tok/sec) represents the average generation speed, while generation time (gen time) shows the average time required for a complete two-step prompt optimization process.
@@ Appendix C Evaluation Statistics
C.1 In-domain test regard to scenery tag
@@
(c) The KDE plot for the Aesthetic Score result of scenery tag test.
(d) The KDE plot for the AI Corrupt Score result of scenery tag test.
Figure 9: The distribution of aesthetic and ai corrupt score for scenery tag test.
The box plot and Kernel Density Estimation (KDE) plot shown in Figure 9 presents the aesthetic scores and AI corruption scores for results from the scenery tag test described in Section 5.2. The analysis clearly demonstrates that TIPO significantly outperforms all other methods, with a substantial margin of improvement.
C.2 In-domain prompt generation test
(a) The box plot for the Aesthetic Score result of short prompt input.
(b) The box plot for the AI Corrupt score result of short prompt input.
(c) The KDE plot for the Aesthetic Score result of short prompt input.
(d) The KDE plot for the AI Corrupt Score result of short prompt input.
Figure 10: The distribution of aesthetic and ai corrupt score for short prompt input.
(a) The box plot for the Aesthetic Score result of truncated long prompt input.
(b) The box plot for the AI Corrupt score result of truncated long prompt input.
(c) The KDE plot for the Aesthetic Score result of truncated long prompt input.
(d) The KDE plot for the AI Corrupt Score result of truncated long prompt input.
Figure 11: The distribution of aesthetic and ai corrupt score for truncated long prompt input.
Figures 10 and 11 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section E.2. While the box plots reveal that the differences in performance between various methods are subtle, the AI corruption scores provide valuable insights. Specifically, these scores indicate that implementations supported by TIPO produce more stable output images compared to other methods.
C.3 Out of focus evaluation
(a) Original Caption
(b) Prompt by GPT4o-mini
(c) Prompt by Prompt-DB
(d) Prompt by Promtist
(e) Prompt by TIPO
Figure 12: The similarity matrix for the 100 best aesthetic results generated in the SD3.5-Large experiments. Off-diagonal elements of the matrix indicate the similarity between different images. A lower value for an off-diagonal element indicates greater diversity among the generated images.
(a) Original Caption
(b) Prompt by GPT4o-mini
(c) Prompt by Prompt-DB
(d) Prompt by Promtist
(e) Prompt by TIPO
Figure 13: The similarity matrix between 100 images of worst aesthetic generated results of SD3.5-Large experiments.
Figures 12 and 13 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  [22]. A matrix with predominantly lower similarity values (overall brighter appearance) indicates high diversity across all generated images, while higher similarity values (overall darker appearance) suggest consistent but less diverse outputs.
For quantitative diversity metrics, please refer to Table 3 in Section 5.
@@ Appendix E Image Examples
E.1 In-domain test regard to scenery tag
@@
E.2 In-domain prompt generation test